Calculate the magnitude and direction of the electric field at a point P which is 30 cm to the right of a point charge Q = -3.0 x 10-6 C


1
Expert's answer
2022-02-18T08:21:57-0500

Given:

"Q = -3.0 *10^{-6}\:\rm C"

"r=0.3\:\rm m"


The magnitude of electric field at the point P

"E=k\frac{|Q|}{r^2}=9*10^9*\frac{3.0 *10^{-6}}{0.3^2}=3*10^5\:\rm V/m"

The field is directed toward the charge.